# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

### v2

@@ Abstract
Instruction-tuned large language models, such as T0, have demonstrated remarkable capabilities in following instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we introduce FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of FarsInstruct dataset coupled with training by Co-CoLA framework [^1] https://github.com/Hojjat-Mokhtarabadi/FarsInstruct, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises more than 200 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability.
@@
FarsInstruct: Empowering Large Language Models
for Persian Instruction Understanding
Hojjat Mokhtarabadithanks: All the authors are members of the Department of Computer Engineering. h.mokhtarabadi@ec.iut.ac.ir Isfahan University of Technology                        Ziba Zamanithanks: Zibazamani@eng.uk.ac.ir Shahid Bahonar University of Kerman
Abbas Maazallahithanks: a.maazallahi@ut.ac.ir University of Tehran                        Hossein Manshaeithanks: manshaei@iut.ac.ir Isfahan University of Technology
@@ 1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs). These models, such as PaLM (Chowdhery et al., 2022), GPT4 (OpenAI et al., 2024), and Llama2 (Touvron et al., 2023) with continuous scaling of their parameters and training data, are known to exhibit emergent properties. Wei et al. (2022a) considers an ability to be emergent if it is not present in smaller models but is present in larger models. This is an unpredictable phenomenon that can not be predicted simply by extrapolating the performance of smaller models. One such ability is instruction-following, which enables models to execute unseen natural language processing (NLP) tasks from reading instructions provided within the input text. Previously, the capability for instruction-following was primarily attributed to the scale of these models. However, recent studies have demonstrated that instruction-following does not exclusively rely on the large size of language models (Sanh et al., 2022). By instruction-tuning on a collection of instructional NLP tasks, smaller language models can learn to follow prompts. This approach has proven to be remarkably efficient, allowing these smaller models to perform competitively and, in some specific tasks, even outperform their larger counterparts (Sanh et al., 2022; Wei et al., 2021; Wang et al., 2022). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning, offering a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
Despite the steady progress of instruction-tuned language models, they still struggle to accurately grasp the subtleties of low-resourced languages due to the scarcity of prompted data and challenges inherent in translating English datasets into other languages (Naous et al., 2024; Ramesh et al., 2023; Vanmassenhove et al., 2021). While efforts have been made to compile extensive multilingual instruction-following datasets, gaps remain in creating diverse and complex prompts for languages like Persian. For example, the SuperNaturalInstructions benchmark (Wang et al., 2022), encompassing various task types across 55 languages, contains merely 2.1% of Persian content. Similarly, the Aya Dataset (Singh et al., 2024), a human-curated effort to enhance AI’s instruction-following abilities across 65 languages, includes 1% of Persian content. This underscores the disparity in the diversity and quantity of the Persian Language tasks compared to other languages.
@@
In this study, we propose FarsInstruct, a comprehensive prompted dataset tailored to the Persian language. It comprises a mixture of manually written instructions ranging from basic to proficient language levels, as well as translations from Public Pool of Prompts (P3) (Sanh et al., 2022) which is a collection of prompted English datasets. In particular, we created more than 200 prompt templates (roughly 10 templates for each of the 21 unique public datasets) that we selected from a variety of sources. These datasets collectively cover ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset, and a detailed overview of FarsInsturct dataset is provided in Section 3.
@@
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in blue, the data field is marked in orange, and the target answer is indicated in gray. In Appendix C, this example is shown in PromptSource environment.
@@
Additionally, in order to facilitate the multi-task adaptation of our model and mitigate the problem of catastrophic forgetting, we introduce Co-CoLA, an integration of CoLA (Xia et al., 2024) with rehearsal training (Kirkpatrick et al., 2017). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we involve retraining a subset of data from previously learned tasks and mixing it with the current task’s data during training. With this periodic revisiting of earlier tasks, the model maintains performance on both old and new tasks while preserving computational efficiency. Section 4 presents an in-depth explanation of Co-CoLA method.
FarsInstruct is publicly available and open-source and we are committed to enhancing it by continually expanding our dataset with a broader range of tasks, instruction entries, and modalities. We hope this dataset fills the critical gap and serves as a valuable resource to the NLP community.
@@ 2 Related work
Instruction-tuning: In the landscape of AI, the capabilities of LLMs have expanded far beyond mere text processing. These sophisticated models are now being fine-tuned in a practice known as instruction-tuning, where models are trained with specific input-output pairs drawn from a wide array of data sources. This technique enables a pre-trained LLM to produce tailored outputs based on given inputs, enhancing its versatility and effectiveness. FLAN (Wei et al., 2021) and T0 (Sanh et al., 2022) pioneered the exploration of instruction-tuned language models, each contributing significantly to the field. FLAN (Wei et al., 2021) adapted a 137-billion parameter pre-trained model, refining it with over 60 NLP datasets using natural language instructions. On the other hand, T0 (Sanh et al., 2022) applied instruction tuning to various T5 models across 2073 prompts from 177 datasets. SuperNaturalInstruction (Wang et al., 2022) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (Muennighoff et al., 2022) expanded on P3’s groundwork (Sanh et al., 2022) by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. In a similar expansive effort, Aya (Singh et al., 2024) emerged as a significant multilingual project, featuring an impressive collection of 513 million instances across 114 languages, achieved through a collaborative research effort that involved fluent speakers from around the world to compile and complete instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging Persian tasks, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
@@
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as the model size and the number of downstream tasks increase. To address this challenge, recent advancements in parameter-efficient fine-tuning methods suggest training only a small portion of parameters while keeping the majority of pre-trained model parameters unchanged. One of the most widely used paradigms in parameter-efficient fine turning is Low-Rank Adaptation (LoRA) (Hu et al., 2021). LoRA only modifies a small, low-rank portion of the model’s weights. This is achieved by adding low-rank matrices to the model’s weights during training. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, and also Kalajdzievski (2024) showed that PEFT strategies, such as LoRA, are still susceptible to catastrophic forgetting. MultiLoRA (Wang et al., 2023) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (Li et al., 2024) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Additionally, CoLA(Xia et al., 2024) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@ 3 FarsInstruct Dataset
With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public prompted dataset for the Persian community. As of this writing, FarsInstruct has more than 200 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more innovative and challenging tasks, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix BC.
@@ 3.1 Dataset Construction
Figure 2: Detailed depiction of 11 task types utilized in our dataset. Each box within the figure lists the specific datasets associated with the respective task type. Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange. Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders. In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
To efficiently create a large collection of prompts, we primarily utilized PromptSource (Bach et al., 2022), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in Bach et al. (2022) is the use of Jinja2 as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Specifically, a template is a function that maps dataset examples into input-output natural language pairs, while a prompt is the combination of an input template and a target template along with a collection of specific meta-data. Instructions are specific directives within input-templates that guide the model’s behavior. However, as the original version of PromptSource did not support Persian, we modified its source code to accommodate Persian datasets. Since this system is originally integrated with Huggingface Datasets (Lhoest et al., 2021) library, we gathered datasets from various sources and consolidated into a unified public repository on HuggingFace [^3] https://huggingface.co/PNLPhub. Appendix C provides some exmaples of prompt template.
@@
In addition to manual templating, we have decided to translate a subset of three question-answering tasks from the P3 dataset Sanh et al. (2022). This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (Costa-jussà et al., 2022) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
Finally, since the datasets were sourced from multiple repositories, we applied a series of pre-processing steps such as deduplication and stripping out non-alphanumeric characters like emojis to ensure normalized text across all data.
3.2 Task Augmentation
@@
It is widely recognized that instruction-tuned models benefit significantly from extensive and varied tasks. Given this context, we focus on developing diverse prompts, spanning from basic to proficient language levels. Furthermore, drawing from the methodologies outlined in FLAN Collection (Longpre et al., 2023), T0 (Sanh et al., 2022), and MetaICL (Min et al., 2022), we enhance task diversity by mixing and swapping different data fields within a given dataset. For instance, whereas a dataset might initially be structured to evaluate a model’s ability to answer question x given input y, we train the model to generate question x when provided with answer y. This approach effectively broadens the spectrum of prompts within a limited data pool.
3.3 Data statistics
The statistics of final dataset after applying templates across all datasets is presented in Figure 3. Table 1 also presents the total number of categorization and generation prompts for each task type.
@@ 3.3 Data statistics
Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes post-prompt application and the number of prompts designed per task type.
@@
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix B.
3.4 Quality Control
We selectively chose publicly available Persian datasets predominantly used for single-task fine-tuning, as their extensive use ensures high quality. Furthermore, to ensure the accuracy and quality of the instructions, we conduct human evaluations through consultations with the general public and experts in the field of literature. This review process allowed us to assess the instructions from multiple perspectives and incorporate cultural and linguistic nuances, critical for ensuring the prompts’ clarity, accuracy, and relevance.

### v3

@@ Abstract
Instruction-tuned large language models have demonstrated remarkable capabilities in following human instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we begin by introducing FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from the Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of the FarsInstruct dataset coupled with training by the Co-CoLA framework, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability. [^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct
@@ 1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (openai2024gpt4), Llama3 (dubey2024llama3herdmodels) and PaLM (chowdhery2022palm). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning and adapts the model’s behavior to respond to user queries with relevant and helpful answers. This technique offers a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
Despite the steady progress of instruction-tuned language models, a persistent limitation remains: their difficulty in capturing the nuanced complexities of low-resource languages. This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages. wang2023far highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation. Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (naous2024having; ramesh-etal-2023-fairness; vanmassenhove-etal-2021-machine). While efforts have been made to compile extensive multilingual instruction datasets (wang-etal-2022-super; singh2024aya; muennighoff2022crosslingual), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets. It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (sanh2022multitask), which is a collection of prompted English datasets. To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets. Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective. For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: ”Can the hypothesis be concluded from the premise? Premise: {Premise}, Hypothesis: {Hypothesis}”, while a corresponding target template could be ”The answer is: {label}”.
@@
The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset after applying its respective template. A detailed overview of the FarsInsturct dataset is provided in Section LABEL:sec:farsinstruct.
@@
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray. In Appendix LABEL:sec:appendix_prompts, this example is shown in the PromptSource environment.
@@
Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (hu2021lora), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (wang2023multilora; li2024mixlora; kalajdzievski2024scaling). To address these issues, we propose Co-CoLA, a novel integration of CoLA (xia2024chain) with rehearsal training (kirkpatrick2017overcoming). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we retrain a subset of data from previously learned tasks, mixing it with the current task’s data during training. This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency. Section LABEL:sec:method presents an in-depth explanation of the Co-CoLA method.
In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation. (2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning. (3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities.
@@ 2 Related work
Instruction-tuning: Instruction tuning refers to the process of training language models using specific input-output pairs derived from diverse data sources. This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language. Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (wang-etal-2022-super; wei2021finetuned), (2) synthesized by prompting proprietary models with a small, seed dataset (alpaca; wang2022self; honovich-etal-2023-unnatural), or (3) generated entirely from scratch, involving human-written prompt-response pairs (DatabricksBlog2023DollyV2; kopf2024openassistant). In this work, we adopt the first approach to develop FarsInstruct. Previous works such as FLAN (wei2021finetuned) and P3 (sanh2022multitask) have been instrumental in advancing instruction dataset creation. FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field. SuperNaturalInstruction (wang-etal-2022-super) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (muennighoff2022crosslingual) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. Similarly, Aya (singh2024aya) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages. This was achieved through a global collaboration involving fluent speakers who contributed instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging tasks in Persian, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
@@
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as model size and the number of downstream tasks increase. To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact. One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (hu2021lora). LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, kalajdzievski2024scaling demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks. MultiLoRA (wang2023multilora) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (li2024mixlora) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Further, CoLA(xia2024chain) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
@@ 3 FarsInstruct Dataset
With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public instruction dataset for the Persian community. As of this writing, FarsInstruct has 197 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more task types, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix LABEL:dataset_detailsLABEL:sec:appendix_dataset_assetsLABEL:sec:appendix_prompts.
@@ 3.1 Dataset Construction
Figure 2: The detailed depiction of 11 task types utilized in our dataset. Each box within the figure lists the specific datasets associated with the respective task type. Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange. Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders. In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
To efficiently create a large collection of prompts, we primarily utilized PromptSource (bach2022promptsource), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in bach2022promptsource is the use of Jinja2333https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Each dataset has multiple prompt templates, each of which consists of an input and a target template. These templates map raw data fields into natural language, structuring both the input and target sequences. Practically, templates allow users to mix arbitrary text with data fields. We refer to the text within the input template that guides the model’s behavior as ”Instruction”. Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
The PromptSource toolkit offers an interface for interactively writing prompts on datasets. However, the original version did not support Persian, so we modified its source code to handle Persian datasets. Our updated version is publicly available, providing the Persian community with a tool to simplify prompt creation  [^4] https://github.com/Hojjat-Mokhtarabadi/promptsource. Appendix LABEL:sec:appendix_prompts depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset. Moreover, since this system was originally integrated with Huggingface Datasets library (lhoest-etal-2021-datasets), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace. Appendix LABEL:sec:appendix_prompts provides a sample of the crafted prompt templates for different datasets.
@@
In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection sanh2022multitask. This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (costa2022no) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
The final dataset is standardized through a series of preprocessing steps like deduplication and removing irrelevant elements (HTML tags, hyperlinks, emojis, and offensive language). Figure 3 shows the distribution of tasks across FarsInstruct, with Table 1 listing the total number of categorization and generation prompts for each task type.
3.2 Task Augmentation and Quality Control
@@
Instruction-tuned models are widely recognized for their significant benefits from exposure to extensive and varied tasks. Given this context, diverse prompts encompassing a range of instructions—spanning from basic to advanced language levels—are essential. Building on best practices outlined in the FLAN Collection (longpre2023flan), T0 (sanh2022multitask), and MetaICL (min-etal-2022-metaicl), we enhance task diversity by mixing and swapping different data fields within a given dataset. For example, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y. This approach effectively broadens the range of prompts available within a limited data pool.
@@
Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes after applying prompt templates and the number of prompts designed per task type. For each dataset, the final size is determined by multiplying the number of samples (N) by the number of prompt templates (M), resulting in a dataset size of N*M.
@@
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix LABEL:sec:appendix_dataset_assets.
To ensure the accuracy and cultural relevance of the instructions, we incorporated public input and expert evaluations. Feedback was gathered from 15 randomly selected individuals and three experts in Persian literature and psychology. Participants were asked to help craft instructions in various writing formats, including formal and informal styles, and to express the same instruction in different ways, then two psychology experts and one literature professor were consulted to refine the instructions. Their expertise informed revisions, ensuring that the responses were grammatically and linguistically correct and resonated with the general Persian-speaking population. Further, the datasets adopted in FarsInstruct are predominantly used for single-task fine-tuning, as their widespread use indicates higher quality.